OpenStudy (littlediva12459):

Determine which pair of points has a negative slope. (–8, 1), (8, –8) (–9, –2), (3, 6) (2, –8), (–1, –9) (6, 9), (5, 7)

563blackghost (563blackghost):

Find the slope of each pair of points....use the slope formula... \(\Huge{slope=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}}\)

563blackghost (563blackghost):

So we will set them up like this... \(\LARGE{A=\frac{-8 - 1}{8- (-1)}}\) \(\LARGE{B=\frac{6-(-2)}{3-(-9)}}\) \(\LARGE{C=\frac{-9 - (-8)}{-1 - 2}}\) \(\LARGE{D=\frac{7-9}{5-6}}\)

OpenStudy (littlediva12459):

I am really bad at this and am really confused on what the answer should be. And would like to thank everyone for the help i just still confused.

OpenStudy (td5):

Follow the given formula @563blackghost gave you. They gave all that is needed to solve the problem. If the answer is negative, then that is your negative slope. A calculator may be ideal for this.

OpenStudy (littlediva12459):

so i got c is this correct

563blackghost (563blackghost):

Not quite. \(\huge {C=\frac{-9 - (-8)}{-1 - 2}} \rightarrow \frac{-1}{-3}=\frac{1}{3}\) Since they are both negative they will cancel out to be a positive so C would not be the answer.

OpenStudy (littlediva12459):

would that make it B because that has a positive and negative number

563blackghost (563blackghost):

\(\huge{B=\frac{6-(-2)}{3-(-9)} \rightarrow \frac{8}{12}}\) Not quite.

OpenStudy (littlediva12459):

Would it be A

563blackghost (563blackghost):

Bingo! ^^

OpenStudy (littlediva12459):

Thank you so much for your help

563blackghost (563blackghost):

Your very welcome ^^

OpenStudy (td5):

If you've not any more questions about how this is solved, please close the question. ^3^